[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Mon Jun 05, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Mon Jun 5 19:52:49 UTC 2017


HSU Shift Log for Mon Jun 05 2017 15:52:44 GMT-0400 (EDT)

Shift: H

Forecaster: M Brennan

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: The ridges workstation was rebooted and did not  
come up properly. Matt Sardi is coordinating with the SDM on addressing  
this.

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170605/4cab9732/attachment-0001.html 


More information about the Ncep.nhc.nco_contacts mailing list